How do you find the slope of (2,3) (7,6)
Answer by askme(9)

the formula for a slope is:
m = (y2 - y1)/(x2 - x1)
this is usually known as "the change in y over the change in x" since you are finding how to get from point A to point B using the coordinates (2,3)(7,6).


x's are on the left of the parentheses and y's are on the right (x,y).


in the coordinates (2,3)(7,6) we will write as this:
y2 = 6
y1 = 3
x2 = 7
x1 = 2


now we substitute these numbers into our formula of a slope.
m = (y2 - y1)/(x2 - x1)
m = (6 - 3)/(7 - 2)
m = 3/5